# SpecDec++: Boosting Speculative Decoding via Adaptive Candidate Lengths

### v2

@@ C.1 Additional Experimental Setups
Datasets. We adopt three datasets in our experiments: (1) Alpaca [39], an instruction-following dataset generated using Self-Instruct [44] from OpenAI’s text-davinci-003 model; (2) HumanEval [9], a test dataset containing Python code synthesis problems; and (3) GSM8K [13], a dataset of high-school math problems. We only use prompts of the datasets and do not use responses.
@@
Inference setting. We set the maximal sequence length to be $512$. We use temperature $T=1$ and adopt top-k sampling with $k=50$. We do not integrate KV cache management techniques such as PagedAttention [22] or KV cache pre-allocation.
@@
C.2 Forward Time Analysis
We report the full results of the linear regression in Section 4.2 in Table 2.
@@ C.2 Forward Time Analysis
C.3 Ablation Studies.
We study how the hyperparameters $w_{\text{rej}},D,h$ influence the final throughputs (tokens per second). First, we calculate the (unweighted) binary KL divergence between the ground-truth probability and the predicted probability, i.e.,
@@
From the table, we see that increasing $w_{\text{rej}}=1$ increases the unweighted eval KL. All the prediction heads trained with $w_{\text{rej}}=1$ perform the best with $h=0.3$ under all three datasets, and similarly, most prediction heads trained with $w_{\text{rej}}=3,6,12$ perform the best with $h=0.5,0.7,0.9$, respectively. This synergy between $w_{\text{rej}}=1$ and $h$ is expected, since increasing $w_{\text{rej}}=1$ forces the acceptance prediction head to focus more on the cases where the candidate token is rejected and thus mitigates the over-confidence issue. In return, the stopping threshold $h$ can be set to a higher value to adjust for the increased predicted probability of existing one rejection.
@@
Appendix D Theoretical Analysis
@@
We analyze the difference $C^{\pi}(s,{\mathrm{continue}})-C^{\pi}(s,{\mathrm{stop}})$ for three cases.
@@
$C^{\pi}(s,{\mathrm{stop}})=c_{1}+c_{2}+C^{\pi}((x_{\text{prefix}}^{\prime},\varnothing)).$
If we choose to ${\mathrm{continue}}$ at the current step, we know that no matter how many additional steps we continue to generate draft tokens, we will eventually discard them and get the same new prefix $x_{\text{prefix}}^{\prime}$. Let $N^{\pi}_{{\mathrm{continue}}}(s)$ be the total number of extra ${\mathrm{continue}}$’s induced by the policy $\pi$ given the current state $s$ and action ${\mathrm{continue}}$. We have
$C^{\pi}(s,{\mathrm{continue}})=c_{1}+c_{1}\cdot(1+N^{\pi}_{{\mathrm{continue}}}(s))+c_{2}+C^{\pi}((x_{\text{prefix}}^{\prime},\varnothing)).$
@@
$C^{\pi}(s,{\mathrm{continue}})-C^{\pi}(s,{\mathrm{stop}})\geq c_{1},\text{ conditioned on }\mathcal{E}_{1}.$
@@
$C^{\pi}(s,{\mathrm{stop}})=c_{2}+C^{\pi}(((x_{\text{prefix}},Y_{1},\dots,Y_{k+1},X_{k+2}),\varnothing)).$
Again, if we choose to ${\mathrm{continue}}$ at the current step, as $Y_{k+2}$ is rejected, future generated tokens beyond $Y_{k+2}$ will also be discarded. After the verification, $Y_{k+2}$ will be replaced by $W_{k+2}\sim\mathrm{Norm}[(p(\cdot|x_{\text{prefix}},Y_{1}\dots,Y_{k+1})-q(\cdot|x_{\text{prefix}},Y_{1}\dots,Y_{k+1}))_{+}]$. Let $N^{\pi}_{{\mathrm{continue}}}(s)$ be the total number of extra ${\mathrm{continue}}$’s induced by the policy $\pi$ given the current state $s$ and action ${\mathrm{continue}}$. We have
$C^{\pi}(s,{\mathrm{continue}})=c_{1}\cdot(1+N^{\pi}_{{\mathrm{continue}}}(s))+c_{2}+C^{\pi}(((x_{\text{prefix}},Y_{1},\dots,Y_{k+1},W_{k+2}),\varnothing)).$
@@
$C^{\pi}(s,{\mathrm{continue}})-C^{\pi}(s,{\mathrm{stop}})\geq c_{1}-\Delta_{1},\text{ conditioned on }\mathcal{E}_{2}.$
@@
$C^{\pi}(s,{\mathrm{stop}})=c_{2}+C^{\pi}(((x_{\text{prefix}},Y_{1},\dots,Y_{k+1},X_{k+2}),\varnothing)).$
If we choose to ${\mathrm{continue}}$ at the current step, there is no immediate cost at the current step and we transit to $(x_{\text{prefix}},(Y_{1},\dots,Y_{k+1}))$.
$C^{\pi}(s,{\mathrm{continue}})=C^{\pi}((x_{\text{prefix}},(Y_{1},\dots,Y_{k+1}))).$
@@
$C^{\pi}(s,{\mathrm{continue}})-C^{\pi}(s,{\mathrm{stop}})\geq-c_{2}-\Delta_{2},\text{ conditioned on }\mathcal{E}_{3}.$
Summary. At the current state, the values of $(Y_{1},\dots,Y_{k})$ are known. We calculate the conditional expectation of $C^{\pi}(s,{\mathrm{continue}})-C^{\pi}(s,{\mathrm{stop}})$ given the current observation. For simplicity of notation, we do not explicitly write out the condition on $(Y_{1},\dots,Y_{k})$.
$\mathbb{E}[C^{\pi}(s,{\mathrm{continue}})-C^{\pi}(s,{\mathrm{stop}})]$
@@
When the right-hand side of the above inequality is larger than zero, the expected total cost of ${\mathrm{continue}}$ is larger than the expected cost of ${\mathrm{stop}}$. Therefore, we obtain a sufficient condition to ${\mathrm{stop}}$ at the current step.

### v3

@@ D.1 Experimental Setups
Datasets. We adopt three datasets in our experiments: (1) Alpaca (Taori et al., 2023), an instruction-following dataset generated using Self-Instruct (Wang et al., 2023) from OpenAI’s text-davinci-003 model; (2) HumanEval (Chen et al., 2021), a test dataset containing Python code synthesis problems; and (3) GSM8K (Cobbe et al., 2021), a dataset of high-school math problems. We only use prompts of the datasets and do not use responses.
@@
Inference setting. We set the maximal sequence length to be $512$. We use temperature $T=1$ and adopt top-k sampling with $k=50$. We do not integrate KV cache management techniques such as PagedAttention (Kwon et al., 2023) or KV cache pre-allocation.
@@
D.2 Forward Time Analysis
We report the full results of the linear regression in Section 5.2 in Table 2. We also visualize $t_{\text{draft}}$ and $t_{\text{target}}$ across the three settings in Figure 4.
@@ D.2 Forward Time Analysis
Figure 4: The forward time of the draft model (llama-2-chat-7B) and the target model (llama-2-chat-70B) under different settings. For each setting, we perform linear regression to calculate the forward times and then average them across different datasets. The additional cost of the acceptance prediction head is negligible compared to the systematic error and the random noise of the environment. Full results are in Table 2.
From Figure 4, we see that the additional cost of the acceptance prediction head is negligible. Besides, in the standalone setting where only the draft model or the target model is used, both $t_{\text{draft}}$ and $t_{\text{target}}$ decrease, which indicates that speculative decoding induces minor overhead in forward passes.
After getting $t_{\text{draft}}$ and $t_{\text{target}}$, we use Equation (5.1) to calculate the theoretical throughputs (tokens per second), which match the noisier empirical throughputs well with relative error $\leq 6.2\%$ for all prompts.
D.3 Ablation Studies.
We fix $w_{\text{acc}}=1$ and study how the hyperparameters $w_{\text{rej}},D,h$ influence the final throughputs (tokens per second). First, we calculate the (unweighted) binary KL divergence between the ground-truth probability and the predicted probability, i.e.,
@@
From the table, we see that increasing $w_{\text{rej}}$ increases the unweighted eval KL. Most of the prediction heads trained with $w_{\text{rej}}=1$ perform the best with $h=0.3$ under all three datasets, and similarly, most prediction heads trained with $w_{\text{rej}}=3,6,12$ perform the best with $h=0.5,0.7,0.9$, respectively. This synergy between $w_{\text{rej}}=1$ and $h$ is expected, since increasing $w_{\text{rej}}=1$ forces the acceptance prediction head to focus more on the cases where the candidate token is rejected and thus mitigates the over-confidence issue. In return, the stopping threshold $h$ can be set to a higher value to adjust for the increased predicted probability of existing one rejection.
@@
D.4 New Experiments on Gemma
We further validate SpecDec++ by repeating all the experiments on a new pair of models: Gemma-1.1-2B-it (Team et al., 2024a) and Gemma-2-27B-it (Team et al., 2024b). The Pareto frontiers of verification rates versus discard rates are plotted in Figure 5, and the empirical speedup is reported in Table 4. Specifically, we see that
  • SpecDec++ has better Pareto frontiers in terms of discard rates v.s. verification rates tradeoff than the baseline SpecDec algorithm on the Alpaca, GSM8K, and HumanEval datasets. The Pareto improvement indicates that SpecDec++ will have better speedups than SpecDec under arbitrary hardware configurations.
  • When deployed on 1 Nvidia A100-80G GPU, SpecDec++ achieves a relative 1.4%, 12.4%, and 7.7% improvement over the baseline methods on the Alpaca, HumanEval, and GSM8K datasets, respectively.
Figure 5: The average verification rates $N_{\text{target}}/N$ and the average discard rates $N_{\text{discarded}}/N$ for SpecDec with different candidate lengths and SpecDec++ with different acceptance prediction heads and stopping thresholds. SpecDec++ has better Pareto frontiers than SpecDec on both the in-distribution dataset Alpaca and the two out-of-distribution datasets HumanEval and GSM8K. The model pair is Gemma-1.1-2B-it/Gemma-2-27B-it.
Table 4: The speedup achieved by SpecDec++ compared to the best speedup achieved by the speculative decoding baseline on Alpaca, HumanEval, and GSM8K datasets. The model pair is Gemma-1.1-2B-it/Gemma-2-27B-it.
Dataset Alpaca HumanEval GSM8K
SpecDec++ 1.35x 2.07x 1.58x
SpecDec (baseline) 1.33x 1.84x 1.47x
relative improvement +1.4% +12.4% +7.7%
We observe fewer improvements on the Alpaca dataset and the GSM8K dataset compared to the llama-2-7b-chat/llama-2-70b-chat model pair. This phenomenon may be caused by weaker alignment between the draft Gemma model and the target Gemma model. On Alpaca, the baseline SpecDec achieves 1.33x speedup for the Gemma model pair but it achieves 1.90x speedup for the Llama model pair; on GSM8K, the baseline SpecDec achieves 1.47x speedup for the Gemma model pair but it achieves 2.07x speedup for the Llama model pair. We see our chosen Gemma model pair performs worse than the Llama model pair, although the model size ratios are roughly the same for the two model families (27B/2B v.s. 70B/7B).
We suspect that for the Gemma-1.1-2B-it and Gemma-2-27B-it model pair on the Alpaca dataset, our acceptance prediction head ends up suggesting an approximately fixed draft length, therefore only achieving a small improvement over the baseline. For settings like math (GSM8K) and coding (HumanEval), the generation involves a mixture of tokens of verbal reasoning and tokens of math calculation/code writing, which naturally require an adaptive draft length.
In summary, although the specific numbers of the improvement vary across different model pairs and different prompt settings, the proposed SpecDec++ indeed achieves an adequate amount of improvements over the baseline. Backed by the theoretical results, our proposed method is a principled extension of the SpecDec method, and adaptively determining the draft length naturally includes the baseline (a fixed draft length) as a special case.
Appendix E Theoretical Analysis
@@
We analyze the difference $C^{\pi}(s,{\texttt{continue}})-C^{\pi}(s,{\texttt{stop}})$ for three cases.
@@
$C^{\pi}(s,{\texttt{stop}})=c_{1}+c_{2}+C^{\pi}((x_{\text{prefix}}^{\prime},\varnothing)).$
If we choose to continue at the current step, we know that no matter how many additional steps we continue to generate draft tokens, we will eventually discard them and get the same new prefix $x_{\text{prefix}}^{\prime}$. Let $N^{\pi}_{{\texttt{continue}}}(s)$ be the total number of extra continue’s induced by the policy $\pi$ given the current state $s$ and action continue. We have
$C^{\pi}(s,{\texttt{continue}})=c_{1}+c_{1}\cdot(1+N^{\pi}_{{\texttt{continue}}}(s))+c_{2}+C^{\pi}((x_{\text{prefix}}^{\prime},\varnothing)).$
@@
$C^{\pi}(s,{\texttt{continue}})-C^{\pi}(s,{\texttt{stop}})\geq c_{1},\text{ conditioned on }\mathcal{E}_{1}.$
@@
$C^{\pi}(s,{\texttt{stop}})=c_{2}+C^{\pi}(((x_{\text{prefix}},Y_{1},\dots,Y_{k+1},X_{k+2}),\varnothing)).$
Again, if we choose to continue at the current step, as $Y_{k+2}$ is rejected, future generated tokens beyond $Y_{k+2}$ will also be discarded. After the verification, $Y_{k+2}$ will be replaced by $W_{k+2}\sim\mathrm{Norm}[(p(\cdot|x_{\text{prefix}},Y_{1}\dots,Y_{k+1})-q(\cdot|x_{\text{prefix}},Y_{1}\dots,Y_{k+1}))_{+}]$. Let $N^{\pi}_{{\texttt{continue}}}(s)$ be the total number of extra continue’s induced by the policy $\pi$ given the current state $s$ and action continue. We have
$C^{\pi}(s,{\texttt{continue}})=c_{1}\cdot(1+N^{\pi}_{{\texttt{continue}}}(s))+c_{2}+C^{\pi}(((x_{\text{prefix}},Y_{1},\dots,Y_{k+1},W_{k+2}),\varnothing)).$
@@
$C^{\pi}(s,{\texttt{continue}})-C^{\pi}(s,{\texttt{stop}})\geq c_{1}-\Delta_{1},\text{ conditioned on }\mathcal{E}_{2}.$
@@
$C^{\pi}(s,{\texttt{stop}})=c_{2}+C^{\pi}(((x_{\text{prefix}},Y_{1},\dots,Y_{k+1},X_{k+2}),\varnothing)).$
If we choose to continue at the current step, there is no immediate cost at the current step and we transit to $(x_{\text{prefix}},(Y_{1},\dots,Y_{k+1}))$.
$C^{\pi}(s,{\texttt{continue}})=C^{\pi}((x_{\text{prefix}},(Y_{1},\dots,Y_{k+1}))).$
@@
$C^{\pi}(s,{\texttt{continue}})-C^{\pi}(s,{\texttt{stop}})\geq-c_{2}-\Delta_{2},\text{ conditioned on }\mathcal{E}_{3}.$
Summary. At the current state, the values of $(Y_{1},\dots,Y_{k})$ are known. We calculate the conditional expectation of $C^{\pi}(s,{\texttt{continue}})-C^{\pi}(s,{\texttt{stop}})$ given the current observation. For simplicity of notation, we do not explicitly write out the condition on $(Y_{1},\dots,Y_{k})$.
$\mathbb{E}[C^{\pi}(s,{\texttt{continue}})-C^{\pi}(s,{\texttt{stop}})]$
@@
When the right-hand side of the above inequality is larger than zero, the expected total cost of continue is larger than the expected cost of stop. Therefore, we obtain a sufficient condition to stop at the current step.